# DESY/PUBDB-2024-01826

### v1

@@ 5 Classifier Performance
To evaluate the performance of DeepJetTransformer, clustered jets from $Z\rightarrow q\bar{q}$ events at $\sqrt{s}=91.2$ GeV and $Z(\to\nu\nu)H(\to q\bar{q})$ events at $\sqrt{s}=240$ GeV were considered. The emphasis was placed on the $Z$ resonance for these studies, with the classification of $H\rightarrow q\bar{q}$ events serving primarily as a comparison to the classification performance of other jet flavour taggers for future colliders, like ParticleNetIDEA Bedeschi_2022 ; Gautam:2022szi . A binary classifier was constructed for each jet flavour $q\equiv u,d,s,c,b,(g)$ with a signal flavour ($i$) and a background flavour ($j$):
@@
where $S_{i}$ are the softmaxed classifier outputs for $Z\rightarrow q\bar{q}$ jets.
@@
Figure 5: The softmaxed classifier scores of the five output nodes of DeepJetTransformer trained with clustered jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV. The contributions of different MC flavours have been displayed.
@@
These classifier scores of the five output nodes of DeepJetTransformer are shown in Figure 5. ROC curves were computed for each $S_{ij}$ combination and are depicted in Figure 6 for the $Z$ resonance and the $ZH$ training. Predictably, the strongest discrimination is between $b$-jets and light jets ($u,d,s$) and is roughly equivalent for all three light jets. The dominant background is from $c$-jets, originating from the similarity of $b$- and $c$-jets with a single reconstructed SV. The discrimination of $c$-jets from the light and $s$-jets likewise cluster together, displaying a similar performance. Figure 6(b) shows that in the low purity regime of $c$-jets, light jets are found to be discriminated worse than $b$-jets before a turnover point at $\epsilon_{sig}^{c}\approx 80\%$, after which distinguishing light jets becomes considerably easier than $b$-jets. It is unclear precisely why this turnover occurs, but it can also be found in ParticleNetIDEA Bedeschi_2022 , and is likely related to $c$-jets with displaced vertices that mimic those of $b$-jets. The sub-leading background comes from $s$-jets, clustered at low to mid charm scores, as also evident in Figure 5(b), possibly due to two reasons: the V0 for some $s$-jets can be misclassified as an SV, and no SVs can be reconstructed for a significant number of $c$-jets.
@@
When $s$-jets are taken to be the signal, as shown in Figure 6(c), $c$- and $ud$-jets present the most challenging backgrounds, with $c$-jets being easier to discriminate against at all signal purities. The $c$-jet background can come from jets where a charm hadron decays to a strange hadron, and only the V0 can be reconstructed, or a strange hadron carries excess momentum. Some discrimination against the dominant $ud$-jets background can be achieved at higher cuts on the strange score, owing to the $K^{\pm}/\pi^{\pm}$ separation and V0 reconstruction. Finally, Figures 5(d) and 5(e) show that classification is most challenging for $u$- and $d$-jets. When $u$-jets are taken to be the signal, it can be seen that DeepJetTransformer learns to discriminate $u$- vs $d$-jets with a $\epsilon_{sig}^{u}\approx 15\%$ at a $\epsilon_{bkg}=10\%$, possibly due to jet charge, which is better than a random classifier, although not considerably.
@@
Figure 6: ROC curves for each $S_{ij}$ combination, as defined in Eq. 4, where $i$ is the signal parton flavour and $j$ is the background flavour. The solid lines correspond to the classification of jets at the $Z$ resonance at $\sqrt{s}=91.2$ GeV, while the dashed lines correspond to the classification of jets from $Z(\to\nu\nu)H(\to q\bar{q})$ events at $\sqrt{s}=240$ GeV.
@@
While considering the performance for $H(\to q\bar{q})$ jets, depicted as dashed lines in Figure 6, no clear trend can be observed. Slight degradation in performance can be observed in the case of $b$ tagging, compared to $Z\rightarrow q\bar{q}$ jets, particularly when $c$-jets are taken to be the background. The discrimination for $c$-jets vs light ($u,d,s$) jets is found to perform relatively the best.
@@
Figure 6(e) shows that the best quark-gluon discrimination can be achieved against the $b$ quarks. This performance can be attributed to several discriminating variables, like jet-constituent multiplicity, constituent momentum distribution, etc., but is likely dominated by the presence or absence of reconstructed SVs. It is the most challenging to discriminate the $s$ and the light quarks from gluons due to their similar jet composition. The prevalent discriminating variable is jet charge, the effect of which is also diluted by the presence of antiquarks and is inferred indirectly by DeepJetTransformer since it is not one of the input variables.
@@
The tagging efficiency of DeepJetTransformer was evaluated for three cases: $b$ vs $c$ tagging, $c$ vs $s$ tagging, $s$ vs $ud$ tagging. Figure 7 shows the efficiency of DeepJetTransformer over the entire jet momentum range and the jet-axis polar angle ($\theta$) range for all three cases. The efficiency for $b$ vs $c$ tagging and $c$ vs $s$ tagging is mostly uniform, showing a good performance for all jet momenta. Similarly, the performance is largely uniform over the $\theta$ range for all three cases, degrading at the extremes due to jet constituents being lost by fiducial cuts.
@@
However, the $s$ vs $ud$ tagging efficiency displays a peculiar distribution over the momentum range of interest, as shown in Figure 7(e). This was found to be dependent on the two most distinguishing features for identifying $s$-jets: $K^{\pm}/\pi^{\pm}$ discrimination and V0 reconstruction. The low-momentum strange jets, on average, have lower $K^{\pm}$ multiplicities, which leads to a reduced tagging efficiency. The very-low-momentum strange jets have a significantly low total charged-particle multiplicity, making V0 reconstruction crucial. The majority of such jets have a single reconstructed V0, making it relatively easier to identify the $s$-jets. On the other hand, the low-momentum strange jets tend to have multiple V0s, splitting the already low jet momentum among these V0s and other hadrons. This is expected to make the strange jet identification more ambiguous. Hence, the $s$-tagging efficiency rises at very low momenta.
5.1 Qualitative Comparison with Other Taggers
A fair quantitative comparison with other taggers developed for future colliders is not feasible due to differing event samples and input features. However, the jet tagging performance trends are very similar to those of ParticleNetIDEA Bedeschi_2022 ; Gautam:2022szi . The strange tagging efficiency of ParticleNetIDEA against the light jets surpasses that of DeepJetTransformer, owing to PID techniques like cluster counting and time-of-flight. A more detailed training dataset is expected to improve the tagging efficiencies of DeepJetTransformer. As one example, performance patterns, indicative of detector design impact, can already be assessed in case of strange tagging with the inclusion of a simple $K^{\pm}/\pi^{\pm}$ classifier while achieving reasonably good tagging efficiencies. The inclusion of V0s facilitates the exploration of another dimension of the detector design through the effects of the vertex and tracking detectors.
DeepJetTransformer outperforms ParticleNetIDEA in bottom-gluon discrimination, especially for efficiencies lower than $90\%$. DeepJetTransformer also has a better discrimination of $b$-jet background for all other signal jet flavours. This efficient discrimination can be attributed to the inclusion of SVs.
With about 106 parameters and efficient transformer blocks as the workhorse, training DeepJetTransformer takes only 2 hours of training to converge after approximately 50 epochs on a NVIDIA Tesla V100s. The computational investment of training is considerably less than the competing architectures, making it an excellent choice to study the constantly evolving detector designs efficiently.
@@ 5.1 Qualitative Comparison with Other Taggers
Figure 7: The jet flavour tagging efficiency with respect to the jet momentum magnitude and the polar angle of the jet axis for jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV. Three cases at $1\%$ background efficiency are shown: $b$ vs $c$ tagging, $c$ vs $s$ tagging, $s$ vs $ud$ tagging.
@@ 5.2 Dependence on the Quality of Particle Identification
The discrimination of $s$-jets is widely regarded as one of the most challenging types of jet discrimination. Thus, it has received considerably less attention than its heavy-flavour counterparts, or indeed gluon discrimination. At the core of the problem is the fact that unlike in the discrimination of quarks vs gluons, which relies heavily on properties following from their differing colour factors $C_{F}=4/3$ vs $C_{A}=3$, or heavy flavour tagging, which relies on displaced vertices of $b/c$ hadrons, strange quarks are treated democratically by QCD and Electroweak theory prior to their decay. Discriminating strange and down jets is particularly challenging due to the same fractional charge of the initiating quarks. In practice, however, strange hadrons carry an excess of the scalar summed momentum of strange jets. This idea was also explored in the context of hadron colliders Nakai:2020kuu . In this work, we exploit the excess momentum carried by strange hadrons, firstly through the inclusion of V0 variables and secondly through $K^{\pm}/\pi^{\pm}$ discrimination.
The $K^{\pm}$ classification scenarios were defined by fixing the efficiency of misidentification to $\pi^{\pm}$ and varying the $K^{\pm}$ identification efficiency. In addition, the limiting cases of Kaon identification with $0\%$ and $100\%$ efficiencies were considered. These are referred to henceforth as the no $K^{\pm}\mathrm{ID}$ and the perfect $K^{\pm}\mathrm{ID}$ scenarios. The considered efficiencies and the misidentification rates are the following:
@@
Figure 8: The dependence of strange jet tagging performance on the inclusion of V0s and charged Kaon identification scenarios.
The largest performance gain with the addition of $K^{\pm}\mathrm{ID}$ information is predictably in the classification of $s$ vs $ud$ jets, shown in Figure 8. Using the no $K^{\pm}\mathrm{ID}$ scenario as a reference, with a $\epsilon_{sig}$ of $31.6$% at a $\epsilon_{bkg}$ of $10$%, strange tagging efficiency improvements of $11.4\%$, $25.9\%$, and $32.9\%$ are evident as the $K^{\pm}\mathrm{ID}$ efficiency is increased to $60\%$, $90\%$, and $95\%$, respectively. The perfect $K^{\pm}\mathrm{ID}$ scenario shows the most sizeable performance gain in $\epsilon_{sig}$ of $82.9\%$. This large performance improvement over the $95\%$ $K^{\pm}\mathrm{ID}$ efficiency with the efficiency of misidentification to $\pi^{\pm}$ of $10\%$ scenario suggests that minimising this misidentification is crucial to strange jet tagging, given their high $\pi^{\pm}$ multiplicity Nakai:2020kuu .
@@
These results indicate the importance and necessity of particle identification techniques, especially for strange quark studies. Such methods are already being explored, like cluster counting and time-of-flight, as foreseen for IDEA CAPUTO2023167969 , and compact-Ring Imaging CHerenkov (RICH) detectors, as being studied for another detector concept for FCC-ee, CLD bacchetta2019cld and ILD, the detector concept developed for ILC https://doi.org/10.48550/arxiv.2203.07535 .
@@ 5.3 Dependence on the Presence of Neutral Kaons
Another distinguishing feature of strange jets is an excess of leading V0s, reconstructed $K_{S}^{0}$ and $\Lambda^{0}$. As noted earlier, these are expected to be more significant in the scarcity of charged Kaons. The inclusion of V0 variables, as Figure 8 shows, results in an improvement of signal efficiency ranging from $16.4\%$ in case of no $K^{\pm}\mathrm{ID}$ to $4.5\%$ in the case of perfect $K^{\pm}\mathrm{ID}$ at a background efficiency of $10$% for $s\text{ vs }ud$ discrimination. This trend proves the importance of V0s to identify strange jets with low $K^{\pm}$ multiplicities or substandard $K^{\pm}/\pi^{\pm}$ discrimination. The performance gain in other forms of classification was again marginal.
A low-material vertex detector with extremely high spatial resolution and a light tracker with numerous measurement points are essential for an accurate track and vertex reconstruction. These, in turn, affect the precise reconstruction and identification of the V0s.

### v4

@@ 4.2.2 DeepJetTransformer Architecture
In summary, the three main differences in the architectures of DeepJetTransformer and ParT (plain) Qu:2022mxj , which both implement a pure Transformer architecture, are the use of attention pooling instead of the typical $cls$ token, additional linear layers prior to the MHA, and the inclusion of the of jet-level variables (listed in Table 2) in addition to jet constituent variables.
@@ 5 Classifier Performance
To evaluate the performance of DeepJetTransformer, clustered jets from $Z\rightarrow q\bar{q}$ events at $\sqrt{s}=91.2$ GeV and $Z(\to\nu\nu)H(\to q\bar{q})$ events at $\sqrt{s}=240$ GeV were considered. The tagger was trained separately for each process. The emphasis was placed on the $Z$ resonance for these studies, with the classification of $H\rightarrow q\bar{q}$ events serving primarily as a comparison to the classification performance of other jet flavour taggers for future colliders, like ParticleNetIDEA Bedeschi_2022 ; Gautam:2022szi . A binary classifier was constructed for each jet flavour $q\equiv u,d,s,c,b,(g)$ with a signal flavour ($i$) and a background flavour ($j$):
@@
where $S_{i}$ are the outputs of the classifier normalised using the SoftMax function, as described in Section 4.2.2. These normalised outputs, which are constrained to lie between 0 and 1 and sum to 1 across all jet flavours, are referred to as ”softmaxed classifier outputs” for brevity.
@@
Figure 5: The softmaxed classifier outputs ($S_{i}$) of the five output nodes of DeepJetTransformer trained with clustered jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV. The contributions of different MC flavours have been displayed.
@@
The five softmaxed classifier outputs of DeepJetTransformer are shown in Figure 5. ROC curves were computed for each $S_{ij}$ combination and are depicted in Figure 6 for the $Z$ resonance and the $ZH$ training. Predictably, the strongest discrimination is between $b$-jets and $s$-, $u$-, $d$- jets and is roughly equivalent for all three background jets. The dominant background is from $c$-jets, originating from the similarity of $b$- and $c$-jets with a single reconstructed SV. Discriminating $c$-jets from $u$-, $d$- and $s$-jets exhibits similar performances, with relatively worse discrimination of the $s$-jet background. Figure 6(b) shows that as the efficiency increases from the right to the left side of the plot, $s$-, $u$- and $d$-jets are discriminated worse than $b$-jets in the high-efficiency regime for $c$-jets until a turnover point at $\epsilon_{sig}^{c}\approx 80\%$, after which distinguishing $s$-, $u$- and $d$-jets becomes considerably easier than $b$-jets. Such a turnover can also be found in ParticleNetIDEA Bedeschi_2022 . The sub-leading background comes from $s$-jets, clustered at low to mid charm scores, as also evident in Figure 5(b), primarily as no SVs can be reconstructed for a significant number of $c$-jets, leaving few variables to distinguish $c$- and $s$-jets.
@@
When $s$-jets are taken to be the signal, as shown in Figure 6(c), $c$- and $ud$-jets present the most challenging backgrounds, with $c$-jets being easier to discriminate against at all signal purities. The $c$-jet background comes from jets where a charm hadron decays to a strange hadron, and only the V0 can be reconstructed, or a strange hadron carries excess momentum. Some discrimination against the dominant $ud$-jets background can be achieved at higher cuts on the strange score, owing to the $K^{\pm}/\pi^{\pm}$ separation and V0 reconstruction. Finally, Figures 5(d) and 5(e) show almost overlapping distributions of classifier scores for $u$- and $d$-jets. Figure 6(d) validates that classification is most challenging for $u$- and $d$-jets. When $u$-jets are taken to be the signal, it can be seen that DeepJetTransformer learns to discriminate $u$- vs $d$-jets with a $\epsilon_{sig}^{u}\approx 15\%$ at a $\epsilon_{bkg}=10\%$, which is better than a random classifier, although not considerably. The discrimination is likely related to a mapping to the initiating parton’s charge, such as the jet charge Field:1977fa ; Krohn:2012fg , the effect of which is diluted by the presence of antiquarks.
(a) bottom tagging
(b) charm tagging
(c) strange tagging
(d) up tagging
(e) gluon tagging
@@
Figure 6: ROC curves for each $S_{ij}$ combination, as defined in Eq. 4, where $i$ is the signal parton flavour and $j$ is the background flavour. The solid lines correspond to the classification of jets at the $Z$ resonance at $\sqrt{s}=91.2$ GeV, while the dashed lines correspond to the classification of jets from $Z(\to\nu\nu)H(\to q\bar{q})$ events at $\sqrt{s}=240$ GeV. The tagger was trained separately for each process. No quark-gluon discrimination results are presented for jets from $Z\to q\bar{q}$ events as the $Z$ boson does not decay into gluons.
@@
While considering the performance for $H(\to q\bar{q})$ jets, depicted as dashed lines in Figure 6, no clear trend can be observed. Slight degradation in performance can be observed in the case of $b$ tagging, compared to $Z\rightarrow q\bar{q}$ jets, particularly when $c$-jets are taken to be the background. The discrimination of $c$-jets vs $s$-, $u$-, and $d$-jets is found to perform relatively the best with respect to the $Z\rightarrow q\bar{q}$ jets when considering the percent-improvement in the ROC Area Under the Curve metric.
@@
Figure 6(e) shows that the best quark-gluon discrimination can be achieved against the $b$ quarks. This performance can be attributed to several discriminating variables, like jet-constituent multiplicity, constituent momentum distribution, etc., but is dominated by the presence or absence of reconstructed SVs. It is the most challenging to discriminate the $s$, $u$, and $d$ quarks from gluons due to their similar jet composition.
@@
The tagging efficiency of DeepJetTransformer was evaluated for three cases: $b$ vs $c$ tagging, $c$ vs $s$ tagging, $s$ vs $ud$ tagging. Figure 7 shows the efficiency of DeepJetTransformer over the entire jet momentum range and the jet-axis polar angle ($\theta$) range for all three cases for two working points. The efficiency for $b$ vs $c$ tagging and $c$ vs $s$ tagging is mostly uniform, showing a good performance for all jet momenta. Similarly, the performance is largely uniform over the $\theta$ range for all three cases, degrading at the extremes due to jet constituents being lost by fiducial cuts.
@@
However, the $s$ vs $ud$ tagging efficiency displays a peculiar distribution over the momentum range of interest, as shown in Figure 7(e). This was found to be dependent on the two most distinguishing features for identifying $s$-jets: $K^{\pm}/\pi^{\pm}$ discrimination and V0 reconstruction. The low-momentum ($3524<|p|<$ 35 GeV) strange jets, on average, have lower $K^{\pm}$ multiplicities, which leads to a reduced tagging efficiency. The very-low-momentum ($|p|<24$ GeV) strange jets have a significantly low total charged-particle multiplicity, making V0 reconstruction crucial. The majority of such jets have a single reconstructed V0, helping identify the $s$-jets. On the other hand, the low-momentum strange jets tend to have multiple V0s, splitting the already low jet momentum among these V0s and other hadrons. This is expected to make the strange jet identification more ambiguous. Hence, the $s$-tagging efficiency rises at very low momenta.
@@
Figure 7: The jet flavour tagging efficiency over the range of jet momentum and the jet axis polar angle for jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV. Three cases at $1\%$ and $10\%$ background efficiencies are shown: $b$ vs $c$ tagging, $c$ vs $s$ tagging, $s$ vs $ud$ tagging.
A similar but exaggerated trend in the distribution is seen for the looser working point of $10\%$ mistag rate for jets with momentum values below 25 GeV. The efficiency is observed to be stable in momentum above this value. As stated above, some of this increase in $s$-tagging efficiency can be attributed to the presence of a reconstructed V0 in jets with low particle multiplicities. Another important aspect to note is that only a small fraction of jets ($<1\%$) with such very low momenta are present in $Z$ boson decays. This means that these low-momentum jets will not have a large contribution to the training of the neural network or the working point determination, which will both be dominated by the bulk of the momentum distribution. The fact that the $10\%$ $u,~{}d$-jet background efficiency also increases to $40\%$ for momenta less than 25 GeV implies that this part of the jet momentum phase space is likely not optimally examined by the neural network. A potential method to improve would be to use training weights flattened over the jet momentum and train on much larger samples with this part of the momentum distribution sufficiently populated. But since these jets contribute to a very small fraction of the total $Z$ boson decays, the improvement in analyses requiring strange tagging would likely not be significant unless the physics case is specific.
5.1 Qualitative Comparison with Other Taggers
A fair quantitative comparison with other taggers developed for future colliders is not feasible due to differing event samples and input features. However, the jet tagging performance trends are very similar to those of ParticleNetIDEA Bedeschi_2022 ; Gautam:2022szi . The strange tagging efficiency of ParticleNetIDEA against the $u$-, $d$-jets surpasses that of DeepJetTransformer, owing to PID techniques like cluster counting and time-of-flight used by ParticleNetIDEA and the conservative PID estimates of DeepJetTransformer. A more detailed training dataset including such PID variables is expected to improve the tagging efficiencies of DeepJetTransformer.
DeepJetTransformer outperforms ParticleNetIDEA in bottom-gluon discrimination, especially for efficiencies lower than $90\%$. DeepJetTransformer also has a better discrimination of $b$-jet background for all other signal quark jet flavours. This efficient discrimination can be attributed to the inclusion of SVs.
With about $10^{6}$ parameters and efficient transformer blocks as the workhorse, training DeepJetTransformer converges within 2 hours after approximately 50 epochs on an NVIDIA Tesla V100s GPU. The computational complexity, measured in FLOPs, is approximately 19.7 MFLOPs. Comparatively, DeepJetTransformer requires fewer FLOPs than competing architectures Qu:2019gqs ; Qu:2022mxj , making it an excellent choice to efficiently test the impact of the constantly evolving detector design on flavour tagging.
@@ 5.2 Dependence on the Quality of Particle Identification
Several $K^{\pm}$ classification scenarios were defined by fixing the efficiency of misidentification to $\pi^{\pm}$ and varying the $K^{\pm}$ identification efficiency. In addition, the limiting cases of Kaon identification with $0\%$ and $100\%$ efficiencies were considered. These are referred to henceforth as the no $K^{\pm}\mathrm{ID}$ and the perfect $K^{\pm}\mathrm{ID}$ scenarios. The considered efficiencies and the misidentification rates are the following:
@@
Table 5: Considered scenarios for $K^{\pm}$ and $\pi^{\pm}$ particle identification performance.
(a) ROC curves
(b) Percent change in signal efficiency
Figure 8: The dependence of strange jet tagging performance on the inclusion of V0s and charged Kaon identification scenarios. (a) ROC curves for $s$ vs $ud$ tagging at the $Z$ resonance at $\sqrt{s}=91.2$ GeV. Solid lines represent results with the inclusion of V0s, while dashed lines show the results without them. (b) Percent change in signal efficiency ($\epsilon_{sig}$) with the inclusion of V0s for $s$ vs $ud$ tagging for each of the $K^{\pm}\mathrm{ID}$ scenarios listed in Table 5. The axes are swapped with respect to Figure 8(a) to present the percent change in signal efficiency ($\epsilon_{sig}$) as a function of 12 fixed background efficiencies ($\epsilon_{bkg}$).
The no $K^{\pm}\mathrm{ID}$ scenario is used as the reference in this section to assess the impact of adding PID variables as input features for jet flavour tagging. The largest performance gain with the addition of $K^{\pm}\mathrm{ID}$ information is predictably in the classification of $s$ vs $ud$ jets, shown in Figure 8. Relative to the reference no $K^{\pm}\mathrm{ID}$ scenario, with a $\epsilon_{sig}$ of $31.6$% at a $\epsilon_{bkg}$ of $10$%, strange tagging efficiency improvements of $11.4\%$, $25.9\%$, and $32.9\%$ are evident as the $K^{\pm}\mathrm{ID}$ efficiency is increased to $60\%$, $90\%$, and $95\%$, respectively. The perfect $K^{\pm}\mathrm{ID}$ scenario shows the most sizeable performance gain in $\epsilon_{sig}$ of $82.9\%$. This large performance improvement over the $95\%$ $K^{\pm}\mathrm{ID}$ efficiency with the efficiency of misidentification to $\pi^{\pm}$ of $10\%$ scenario suggests that minimising this misidentification is crucial to tagging strange jets, given their high $\pi^{\pm}$ multiplicity Nakai:2020kuu .
@@
These results confirm the importance and necessity of particle identification techniques, especially for strange quark studies, as was also noted by some previous studies Azzi2021 ; Albert:2022mpk ; Bedeschi_2022 .
@@ 5.3 Dependence on the Presence of Neutral Kaons
As noted earlier, an excess of V0s, reconstructed $K_{S}^{0}$ and $\Lambda^{0}$, carrying the bulk of the jet momenta is also a distinguishing feature of strange jets and these are expected to be more significant in the scarcity of charged Kaons. The inclusion of V0 variables, as Figure 8 shows, results in an improvement of signal efficiency ranging from $14.3\%$ in case of no $K^{\pm}\mathrm{ID}$ to $4.2\%$ in the case of perfect $K^{\pm}\mathrm{ID}$ at a background efficiency of $10$% for $s\text{ vs }ud$ discrimination. The percent improvement in signal efficiency for each of the $K^{\pm}\mathrm{ID}$ scenarios listed in Table 5 is depicted separately in Figure 8(b). This trend proves the importance of V0s to identify strange jets with low $K^{\pm}$ multiplicities or substandard $K^{\pm}/\pi^{\pm}$ discrimination. The performance gain in other forms of classification was again marginal.